Journalist Report – March 13th

Sol 17

"Emergency return"

When the EVA team of the day was exploring the South area of the station
they faced strong winds and took the decision to get back to the
station. The beginning of the EVA rolled out fine : they went to the
weather station to do some maintenance, as usual. Then they headed
south. The first stop was done before the Kissing Camels Ridges. The
wind was still low enough, but it was getting colder. It’s at their
second stop that things started to get serious : after 10 minutes
walking, the wind increased dramatically and quickly, the EVA team was
starting to feel the cold of the wind through their gloves. It was
getting hard to hear other members through the radio, and the wind was
pushing them around. Cerise, EVA leader, took the right decision to
abort the EVA and head straight back to the rovers. The team drove right
back to the Hab. The wind strained to rover’s batteries quite a lot. The
team arrived in the Hab feeling frozen, but safe and sound.
This EVA was different from the rest of our mission, Cerise, Jérémy,
Aurélien and Norbert explored a wide region without a precise goal
except the will to explore and discover unknown areas. They drove the
rovers along Cow Dung Road and stopped each time they found a nice
geologic spot. That’s real exploration, like the first settlers will
have to do on Mars.

Today was a milestone for some of us, our mission will end in less than
two days and we started to realize that today. The effect on our
behaviors is strange : we have a foot on Earth but the other one still
sealed on Mars. On one hand it’s easier for us to live in this
philosophy as confinement has less weight on us. But on the other hand
the mission is one of the most intense moments of our lives that we
prepared for one year and it’s hard to see its end coming. Two of us
will return to the MDRS to be commander of the 2020 missions of
ISAE-SUPAERO but for the others these are certainly our last days on the
Red Planet. It affects the life of the Hab and the experiments. But it’s
very different from a crew member to another. For example Aurélien and
Norbert felt the end of the mission coming very fast while I’m still
feeling like an astronaut in the middle of its rotation, I was really
surprised when Aurélien told me that tomorrow we will start to pack our
luggage. The fact that in three days we will be on a plane to Paris
makes us feel different in these days than if we were in the middle of a
longer mission. We are thinking about the return more than ever and that
affects us, and I think that these days, like the first days of the
mission are less pertinent to the extent of science and human factors.
On Saturday, crew 207 will join us in the MDRS. The Hab and the crew
will be ready to welcome them and let them start science here.

Operations Report – March 13th

Crew 206 Operations Report 13 – MAR – 2019

SOL: 17

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and
alternator need replacement on generator. A contractor has
been arranged to make these repairs within a few days.

Generator (hours run): 13,5 hours

Solar— SOC % – Turned off at 100%, Still off around 90% (1930)

Diesel Reading – 60%
Propane Reading – 73%

Ethanol Free Gasoline – 2 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – 330 gallons

Water differential (static tank) – 10 gallons

Gallons transferred: Not Applicable

Water in GreenHab – 208 gallons

Static to Loft Pump used – Yes

Water Meter: 0141756,4 gallons

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Breaking system being repared.

Hours: 53,1 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: No

Curiosity rover used: Used

Hours: 81,0 hours

Beginning charge: 100%

Ending charge: 74%

Currently charging: Yes

Spirit rover used: Used

Hours: 82,4 hours

Beginning charge: 100%

Ending charge: 55%

Currently charging: Yes

Notes on rovers: Opportunity’s brake fluid is empty, it looks like the
pipeline has a leak. Atila is working on it, more informations will follow.
On today’s EVA, Curiosity had another issue with its
direction. We are working on it with Atila.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Not Used.

Staff Crew Car used and why, where? Off Campus

General notes and comments: Nothing to report.

Summary of internet: Nothing to report.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing more to report.

Greenhab Report – March 13th

Crew 206 Green Hab Report – 13-MAR-2019

GreenHab Officer: Norbert Pouzin

Environmental control: Heating and Cooling w/ ambient air (0 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 17.3°C

High: 26.5°C

Hours of supplemental light: 0 Hours

Daily water usage for crops: 4.2 gallons

Water in Blue Tank: 208.6 gallons

Crops watered at 0820 hours (2.1 gallons) and 1840 hours (2.1 gallons):


1) 08.20 AM 24.1° C / 25 %
2) 11.30 AM 22.3° C / 20 %
3) 15.00 PM 22.8° C / 20 %
4) 18.40 PM 24.1° C / 20 %

Changes to crops: Lettuces, savoury mix, and mixed green salads planted
this week are sprouting.

It was a cold day today, so the heater worked almost all day long.
We harvested herbs and carrots for our diner.

Harvest: 88 g of carrots, 2 g of cress, 2 g of parsley and 3 g of sage

Support/supplies needed: None.

EVA Report – March 13th

Crew 206 EVA Report 13Mar2019

EVA #13

Author: Cerise Cuny

Purpose of EVA: This EVA will allow us to do some maintenance on the
LOAC and weather station (hygrometry temperature captor). In addition to
that, we will go on an exploration hike at 518600/4248500 that today’s
EVA teams spotted as an interesting place.

Start time: 9:14 am
End time: 10:42 am

Narrative: After the 5 minute depressurization, we checked the water
levels and the rovers’ batteries. We left the station at 9:26 am and
arrived to the LOAC three minutes later. We placed the hygrometry and
temperature captor at the feet of the LOAC, in the same plastic bag than
the battery to protect it from the rain. We left the LOAC location at
9:29, lost communication with the hab at 9:32 am. We did a first stop
not far before Kissing Camels Ridges, and walked a bit (518214/4249859).
At 9:55 we took back the rover to go farthest to the south. We stopped
at 10:03 am at 519433/4248119. We walked for a bit in canyons but
quickly the wind became stronger. I decided to turn around, to go back
directly to the rovers and to the hab. At some point, we couldn’t hear
very well what people were saying in the radios, and we were pushed by
the wind. We arrived to the rovers at 10:19 am and after an 11 minutes
drive we arrived to the station. We picked up some garbage bags that
were pushed by the wind. We entered the main airlock at 10:37, grateful
to be protected from the cold wind.

Destination: LOAC and Weather Station location, Points of interests in
the south.

Coordinates (use UTM NAD27 CONUS): 518500/4257000 (LOAC and Weather
Station), 518214/4249859, 519433/4248119.

Participants: Jérémy Auclair (CMD), Aurélien Mure (XO), Cerise Cuny
(HSO, EVA leader), Norbert Pouzin (GHO)

Road(s) and routes per MDRS Map : Entrance Road , Cow Dung Road

Mode of travel: Driving and Walking

Sol Summary – March 13th

Crew 206 March 13th 2019
Sol 17

Summary Title : Last days on Mars

Author : Jeremy Auclair

Mission Status : All nominal

Sol Activity Summary : In two days we are going to break the simulation
and wrap up all our experiments, it will be the end of a great
adventure. We are feeling this, everything is going faster, days are
flying by. We are savouring these last days. This morning went fine,
although we had to shorten our EVA. While we were exploring, and after
the weather station maintenance, the wind started increasing greatly, to
the point where we hard a hard time hearing ourselves in the radio and
we were being pushed around. We headed straight back to the rovers and
drove back to the Hab. Nothing bad happened. The wind has been blowing
strong since then. The weather station is still up. This afternoon was
quite busy : the TELEOP session went fine and due to planning
constraints we all had to watch VR videos, which took some time. All our
experiments are going fine, although the crew is being less attentive to
water consumption measurement.

Look ahead plan : Tomorrow’s EVA will be used to test the prototype
space suit again.

Anomalies in work : None

Weather : Cloudy all day, very windy since 10:00 am

Crew physical status : Fine

EVA : Weather station maintenance and South MDRS exploration

Reports to be filed : GreenHab, Journalist, Operation, EVA Request, EVA

Support Requested : MDRS weather station data access.

Copyright © The Mars Society. All rights reserved. | Main Site